If f(x) = sqrt(x) and g(x) = x - 2 find (f og)(27)
Answer by stanbon(75887)

If f(x) = sqrt(x) and g(x) = x - 2 find (f og)(27)
----
fog(27) = f[g(27)]
= f[27-2]
= f(25)
= sqrt(25)
= 5
--------------
Cheers,
Stan H.
================